OpenStudy (anonymous):

Write a direct variation equation that relates x to y. Then solve. Show both the equation and the solution. If y = 8 when x = 4, find y when x = 5

OpenStudy (skullpatrol):

Is this exactly how the question is worded in your assignment?

OpenStudy (anonymous):

Yea.

OpenStudy (skullpatrol):



OpenStudy (anonymous):

multiply it by two? or add four?

OpenStudy (skullpatrol):

Yes, either will work but the question asks you to write a "direct variation" equation that relates x to y. Which means to multiply by 2. Can you write the equation?

OpenStudy (anonymous):

4x*8 = 5x*y?? I'm not good with finding equations

OpenStudy (skullpatrol):

That is way too complicated, what you want is something simpler in the form: y = k*x, where k is a constant.

OpenStudy (anonymous):

so y=2*x or 8=2*4?

OpenStudy (skullpatrol):

Correct.

OpenStudy (skullpatrol):

Now find y when x = 5?

OpenStudy (anonymous):

10?

OpenStudy (skullpatrol):

Correct.

OpenStudy (anonymous):

Thanks!

OpenStudy (skullpatrol):

np